Circular No. 6132

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995B IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of an apparent supernova on Technical Pan films (limiting mag ∼ 21.5) obtained on Jan. 6.00 and 9.06 UT with the OCA 0.9-m Schmidt telescope. Then at magnitude V ∼ 19.0, SN 1995B is located at α = 7^{h}17^{m}33^{s}.82, δ = +21^{o}49′25′′.0 (equinox 2000.0), which is 2′′.8 west and 7′′.3 south of the galaxy’s center. On Feb. 7.0, the supernova is still present but ∼ 1 mag fainter than in early January. A nearby star of mag 16 has position end figures 34^{s}.72, 58′′.6. The limiting magnitude of the discovery film is ∼ 21.5. Nothing is visible at the location of SN 1995B on films of limiting mag 22 obtained at OCA in 1989 Dec. and on 1990 Nov. 26.1.

KR AURIGAE K. Honeycutt and J. Robertson, Indiana University (IU), write: “KR

Aur occasionally falls from its high state of V = 12–13 to 17–18. The IU Roboscope shows no low states during 1991–94 (347 observations), but over the last 100 days, KR Aur has faded smoothly to V = 16.5 and is still falling.”

SUPERNOVA 1993J in NGC 3031 A. P. S. Crotts, Columbia University, reports: “SN 1993J is accom-

panied by a companion at radius 0′′.84 ± 0′′.09 and position angle 344^{o} ± 9^{o} from the supernova, consistent with the north-south elongation seen in pre-explosion images of the progenitor (IAUC 5758). The data were obtained at the Canada-France-Hawaii Telescope on 1994 Dec. 30 UT in two medium-width bands at 684 and 819 nm, with the FWHM of images as good as 0′′.60. No structure besides the two point sources is seen. The magnitude of the companion, based on these bands, is I ∼ 22.7 (R − I ∼ 0.5; uncertainties of ∼ 0.2 mag). The companion will affect late-time photometry of the supernova.”

NOVA CIRCINI 1995 Visual magnitude estimates: Feb. 1.934 UT, 8.6 (T. Cooper, Benoni,

S. Africa); 2.063, 8.1 (D. Overbeek, Edenvale, S. Africa); 3.853, 8.9 (L. A. G. Monard, Pretoria, S. Africa); 4.951, 8.8 (Monard).

1995 February 7 Daniel W. E. Green